PLTOTF(1) PLTOTF(1) NAME pltotf−convertpropertylistfilestoTeXfontmetric(tfm)format SYNOPSIS pltotf[−verbose]plfilenametfmfilename DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thepltotfprogramtranslatesa(human-oriented)WEBpropertylistfiletoa(program-oriented) TEXfontmetricfile.Thus,aftereditingthepropertylistfile,aTFMfilecanbegeneratedforuse with,forexample,tex(1). Boththeplfilenameandthetfmfilenamemustbecomplete;noaddingofdefaultextensionsor pathsearchingisdone. OPTIONS Withoutthe−verboseoption,pltotfoperatessilently.Withit,abannerandprogressreport areprintedonstdout. SEEALSO tftopl(1). DonaldE.Knuth,TEXware. DonaldE.Knuth,TheMETAFONTbook,(VolumeCofComputersandTypesetting),Addison-Wes-ley,1986,ISBN0-201-13445-4. AUTHORS DonaldE.Knuthwrotetheprogram.ItwaspublishedaspartoftheTEXwaretechnicalreport, availablefromtheTEXUsersGroup.HowardTrickey,PavelCurtis,andRichardFurutaoriginallyportedittoUnix. Web2C7.32x 27December1992 1